OpenStudy (anonymous):

(double integral)R 1/(1+x^2+y^2)^2; R is the first quadrant. Solve the integral

OpenStudy (turingtest):

\[\int\int{1\over1+x^2+y^2}dA\]in the first quadrant?

OpenStudy (turingtest):

if so, convert to polar coordinates

OpenStudy (anonymous):

1/(1+x^2+y^2)^2

OpenStudy (turingtest):

right sorry\[\int\int{1\over(1+x^2+y^2)^2}dA\]let's convert these to polar coordinates and see what happens...

OpenStudy (anonymous):

how do you do that?

OpenStudy (turingtest):

for polar coordinates\[x=r\cos\theta\]\[y=r\sin\theta\]\[dA=rdrd\theta\]

OpenStudy (turingtest):

isn't there a given region?

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

there is but its long wait i am gonna try to write it

OpenStudy (turingtest):

like integrate over the first Q of x^2+y^2=1 or something?

OpenStudy (turingtest):

well we need the region to find the bounds on r I think

OpenStudy (anonymous):

limit b to infinity, (integral)alfa to beta (integral) a to b g(r,theta) r dr dtheta

OpenStudy (anonymous):

this is what it says in the book

OpenStudy (turingtest):

that is just a general formula for what I am showing you nevermind, just do the polar conversion I mentioned earlier

OpenStudy (anonymous):

oh okay lol

OpenStudy (anonymous):

how do i go about solving it

OpenStudy (turingtest):

do the substitution I mentioned earlier and you will see the integration is easy I just don't know what the bounds on r are supposed to be

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i don't even know how to set it up

OpenStudy (turingtest):

yeah, it looks like the bounds on r are from 0 to 1 so just plug in the subs for conversion to polar coordinates I gave above and simplify\[x=r\cos\theta\]\[y=r\sin\theta\]\[dA=rdrd\theta\]plug that into your integral\[\int\int{1\over(1+x^2+y^2)^2}dA\]

OpenStudy (anonymous):

(double integral) 1/(1+(r cos theta)^2 + (r sin theta)^2)^2 r dr dtheta ?

OpenStudy (turingtest):

yes, now simplify that using trig identities hint: factor out r^2 where possible, that should reveal a trig identity

OpenStudy (anonymous):

costheta + sintehta is cos2theta right?

OpenStudy (turingtest):

I don't think so, but that's not the identity I had in mind let's look at just what is in the parentheses in the denom...

OpenStudy (turingtest):

\[1+(r\cos\theta)^2+(r\sin\theta)^2\]simplify this using the hint I gave you

OpenStudy (anonymous):

hmm....

OpenStudy (turingtest):

I'll let you think about it for a minute while I help someone else :) hint: MAJOR trig identity here after factoring out r^2

OpenStudy (anonymous):

1+r^2? lol sorry i suck at math

OpenStudy (turingtest):

that is right! and how did you make it to multivariable calc if you suck at math?

OpenStudy (anonymous):

i mean i don't suck but i like blank out sometimes

OpenStudy (turingtest):

anyway, our integral is now\[\int\int{1\over(1+r^2)^2}rdrd\theta\]

OpenStudy (turingtest):

we are in Q1, so what are the bounds on theta?

OpenStudy (anonymous):

0 to 2pi?

OpenStudy (turingtest):

that would be a full circle (we would go through all 4 quadrants that way) we want just Q1 so think again

OpenStudy (anonymous):

hmm. 0 to pi

OpenStudy (anonymous):

oh wait its 0 to pi/2 right?

OpenStudy (turingtest):

yes :)

OpenStudy (turingtest):

now the only thing I am unsure of is the limits of r, since no region is given I think it may be from 0 to 1 but am not sure... there is no equation for the region R in your problem?

OpenStudy (anonymous):

no not in this problem but in problem before its 0 to infinity and 1 to infinity. but in this problem its not given

OpenStudy (turingtest):

let me check my notes...

OpenStudy (turingtest):

my notes are not helping me the phrasing is strange for me, but since Q1 is an infinite region making the bounds on r 0 to infinity is plausible never seen it before though, but it fits with your notes it seems

OpenStudy (anonymous):

ye i don't know why i gives limits for r in this problem. its there in all the other problem

OpenStudy (turingtest):

it does \(not\) give the limits for r, right?

OpenStudy (turingtest):

if not, then the integral wil then be\[\int_{0}^{\frac\pi2}\int_{0}^{\infty}{1\over(1+r^2)^2}rdrd\theta\]if they are given, use them instead gotta go, good luck!

OpenStudy (anonymous):

thanks